OpenStudy (dancing_stars23):

Find the sixth term of a geometric sequence with t5 = 24 and t8 = 3.

iYuko (iyuko):

t5 = a*r^4 = 24 t8 = a.r^7 = 3 So now what do you do?

OpenStudy (dancing_stars23):

I just saw the exact same post and to me that didn't make any sense

OpenStudy (dancing_stars23):

@♂

iYuko (iyuko):

We are dividing one by the other... \[R ^{3}\] = 1/8 r = 1/2 t6 = a.r^5 = t5*r = 24 * 1/2 = 12

iYuko (iyuko):

R^3=1/8

OpenStudy (dancing_stars23):

the first part makes sense and the second but how did you get r^3 = 1/8 That's the part I don't get. I know you said divide, and 24/3 equals 8. but why 1/8

OpenStudy (anonymous):

the way I'd do it, is: \(24=t_5\\24*r=t_6\\ 24*r^2=t_7\\ 24*r^3=3=t_8\\24*r^3=8\implies r^3=\dfrac{1}{8}\\ \) solve for r

iYuko (iyuko):

r=1/2

OpenStudy (dancing_stars23):

@♂ thanks! that makes more sense now

OpenStudy (xapproachesinfinity):

hmm the geometric sequence's Nth term is \[a_n=ar^{n-1}\] where a is the first term and r is the common ratio so our goal here is to find those two given the info

OpenStudy (xapproachesinfinity):

\[t_8=tr^{7}=3 \\ t_5=tr^4=25\]

OpenStudy (xapproachesinfinity):

from those two equations we can find what t the first term and r the common ratio are

OpenStudy (xapproachesinfinity):

then we can find the sixth term or any other term we desire to find

OpenStudy (xapproachesinfinity):

tr^4=24 not 25 correction! and 1/2 is correct as suggested by the fellows above :)

OpenStudy (xapproachesinfinity):

well didn't see @iYuko work decent work as well grant him/her a medal too :)